OpenStudy (rhr12):

HELP PLEASE:-------- 1/4(4/3x + 16) =2(x+3)

OpenStudy (rhr12):

@pooja195 @mathstudent55

OpenStudy (anonymous):

If I were you I would distribute first

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{4}\left( \cfrac{4}{3x+16} \right)=2(x+3)?\)

OpenStudy (anonymous):

he probably means \[\frac{ 1 }{ 4 }( \frac{ 4 }{ 3 }x + 16) = 2(x+3) \]

OpenStudy (rhr12):

@nonon is right and i already got the answer. It's -6/5